OpenStudy (anonymous):

Which of the following is the equation of a circle whose center is at the orgin and whose radius is square root 5? 1. x2+y2=squareroot 5 2. x2+y2=5 3. x2+y2=10 4. x2+y2=25

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[x ^{2}+y^2 = r^2\]

OpenStudy (anonymous):

where r is the radius

OpenStudy (anonymous):

so u have to find equation whoz radius is root 5

OpenStudy (anonymous):

\[so r^2 =(\sqrt{5})^2 =5\]

OpenStudy (anonymous):

so .......r^2 **

OpenStudy (anonymous):

\[r^2 = (\sqrt{5})^2 =5\]

OpenStudy (anonymous):

so you get x2+y2=5

OpenStudy (anonymous):

Thanks you alot Been stuck on it for a while

OpenStudy (anonymous):

you are welcome